OpenStudy (anonymous):

Find the value of the variable(s). if you're answer is not an integer , leave it in simple radical form. The triangle is not drawn to scale.

OpenStudy (anonymous):

\[a) x=17, y=34\sqrt{3}\] \[b) x=34, y=17\sqrt{3}\] \[c) x=34\sqrt{3}, y=17\] \[d) x=17\sqrt{3}, y=34\]

OpenStudy (anonymous):

@jigglypuff314

jigglypuff314 (jigglypuff314):



OpenStudy (anonymous):